# Convolutional Neural Network NLP Natural Language Processing

### v1

@@ Abstract
The generalization of the Transformer architecture via MetaFormer has reshaped our understanding of its success in computer vision. By replacing self-attention with simpler token mixers, MetaFormer provides strong baselines for vision tasks. However, while extensively studied on natural image datasets, its use in medical imaging remains scarce, and existing works rarely compare different token mixers, potentially overlooking more suitable designs choices. In this work, we present the first comprehensive study of token mixers for medical imaging. We systematically analyze pooling-, convolution-, and attention-based token mixers within the MetaFormer architecture on image classification (global prediction task) and semantic segmentation (dense prediction task). Our evaluation spans eight datasets covering diverse modalities and common challenges in the medical domain. Given the prevalence of pretraining from natural images to mitigate medical data scarcity, we also examine transferring pretrained weights to new token mixers. Our results show that, for classification, low-complexity token mixers (e.g. grouped convolution or pooling) are sufficient, aligning with findings on natural images. Pretrained weights remain useful despite the domain gap introduced by the new token mixer. For segmentation, we find that the local inductive bias of convolutional token mixers is essential. Grouped convolutions emerge as the preferred choice, as they reduce runtime and parameter count compared to standard convolutions, while the MetaFormer’s channel-MLPs already provide the necessary cross-channel interactions. Our code is available on GitHub.
@@ 1 Introduction
Over the last ten years, Deep Learning and especially Convolutional Neural Networks have set the state-of-the-art in various tasks of medical image processing like classification [1], segmentation [2] and detection [3]. Following its success in Natural Language Processing (NLP[4, 5], self-attention [6] has soon emerged as an alternative to convolution. The Transformer architecture has been successfully adapted to image classification [7], segmentation [8] and detection [9]. Further adaptation of the Transformer architecture targeted specific requirements for vision tasks. This includes for example a hierarchical pyramid-like architecture like it can be found in CNNs [10, 11], the limitation of its receptive field to the direct neighbourhood for detection [12], or tackling characteristics of self-attention that hinder the usage of Vision Transformers as the backbone for further vision tasks [13].
@@
The property of self-attention having a global receptive field is a doubled-edged sword in image processing where large input sizes are common. On the one hand, the number of trainable parameters of the self-attention operator is independent of the input size, and it enables capturing of large context. On the other hand, since every pixel attends to each other pixel in an input, its computational complexity grows quadratically with the resolution of 2D images, making it inefficient for high resolutions. In contrast, convolution limits its receptive field to the direct neighbourhood of the pixel and thus provides a strong local bias and translation-equivariance, two properties fitting the domain of images. Combining both paradigms, many works have also limited the receptive field of self-attention, introducing local self-attention. Here, the implementations can be divided into two groups, splitting the input into patches[14] or limiting the influence of pixels outside the direct neighbourhood by applying a (soft) weighting during the self-attention operation [15]. For both local and global attention, implementations were introduced [16, 17]. Further, recent work investigates into the principle of filter separation, bringing state space models to computer vision within the MAMBA architecture [18]. Lately, frameworks like PyTorch’s flexattention enables the easy implementation of various attention patterns.
@@
There also exist efforts (e.g. classification [19] and segmentation [20, 21]) to enable large receptive fields for CNNs without equipping the model with too many learnable parameters.
@@
While [22] demonstrates that replacing the self-attention within the ViT with Multi-Layer Perceptron (MLP) yields comparable performance [7, 14], the MetaFormer framework [23] takes this a step further by abstracting the general transformer architecture into two parts: The token mixer (e.g. self-attention) learns to aggregate spatial information and the channel-MLP extracts new features with a point-wise operation. Within this MetaFormer perspective, using grouped convolution as the token mixer naturally draws a parallel to ConvNeXt [19], where depthwise convolution fulfils an analogous role in aggregating spatial information. However, using pooling as a simple, non-parameterised token mixer, the MetaFormer outperforms ViT on ImageNet. This implies that the channel-MLPs provide the main contribution to the transformer’s capabilities in computer vision, rather than the token mixer. In their following work, Yu et al. [24] provide strong results (ImageNet top-1 accuracy of over 80%) using no token mixing at all. Still, equipping the token mixers with some learnable parameters was found beneficial (see [23, Tab. 5]).
@@
The general formulation of the MetaFormer easily allows for the deployment of hybrid architectures, combining different token mixers (e.g. using pooling as an efficient token mixer in the first stages with high-resolution images and global self-attention in latter ones). Since then, MetaFormer has been implemented in various applications and extended with specialised token mixers (see Sec. 1.2).
@@ 1.1 Contribution
Most existing applications of the MetaFormer lie outside the medical domain, and those within typically employ it without comparing different token mixers – potentially overlooking more suitable choices. This work presents the first comprehensive study of token mixers for medical imaging, analysing their strengths, weaknesses, and complexity in terms of trainable parameters and runtime. We systematically compare six token mixers, varying the kernel size for kernel-based mixers such as convolution and pooling. Our evaluation covers eight datasets, including seven from the medical domain and one natural image benchmark. We investigate two different tasks: classification as a global prediction task (five datasets) and semantic segmentation as a dense prediction task (three datasets, covering small, medium, and large resolutions). The datasets represent diverse modalities, including histology, microscopy, and radiography. With this, we chose to expose the different token mixers to key medical imaging challenges such as large homogeneous regions in abdominal CT and long-range context requirements in pathology. We note our key findings as followed:
  • Datasets containing a single dominant object in the image favour more locality and thus smaller kernel sizes. Vice versa, datasets with a larger region of interest tend to profit from incorporating a broader spatial context through larger kernels.
  • For classification, MetaFormer trained from scratch yielding the best performance when using simple token mixer.
  • When pretrained weights available, exchanging token mixers introduce a domain gap.
  • Pretrained weights of global attention can be successfully utilized in local attention. Providing a more efficient alternative when dealing with large input resolutions.
  • No token mixing already yields competitive classification performance.
@@
1.2 Related Work
With the right choice of token mixer, the MetaFormer can be a light-weighted model, enabling its use for real-time applications like ship signature classification in Synthetic Aperture Radar (SAR) [25] or in autonomous driving [26], which deploys additional token pruning for reduced runtime complexity. For super resolution of meteorological satellite cloud images, [27] introduces a new token mixer combining gated convolution and channel-space attention. With the extension of temporal pooling, the MetaFormer has been applied on time series, capturing the structural and temporal relationship in human motion prediction for human-robot interaction [28]. A novel idea of token mixing moves this operation to the Fourier spectrum, enabling efficient global token mixing without limited receptive field, but adding the cost of the Fourier transformation and its inverse. With balancing the low- and high-pass characteristics via spectral pooling aggregation [29], the performance on ImageNet is improved by 3% accuracy compared to the PoolFormer (MetaFormer with average pooling as token mixer) [23]. In [30], this approach is extended, so coefficients of a dynamic filter bank are estimated by an MLP based on the input. While the original work [23] employs the MetaFormer itself as a backbone for semantic segmentation, MetaSeg [31] extends its principles to a decoder and introduces channel reduction attention for gathering global context within the decoder.
In the medical domain, MetaFormer mostly gets adapted for semantic segmentation. For polyp segmentation, MetaFormer has been paired with a parallel path comprising convolutional units with residual and dilated paths to compensate for MetaFormer’s weak low-level features [32]. Another work on polyp segmentation [33] utilizes the MetaFormer as the encoder in a U-Net-like architecture and extend the decoder’s dilated convolutions with learnable multiscale upsampling. In volumetric brain tumour segmentation, MetaFormer has been used as an auxiliary decoder, combining local–global pooling with depthwise convolution as the token mixer. To improve efficiency, it has been adapted with involution and a squeeze-and-excitation module in place of channel-MLP [34]. Similarly, MAMBA’s structured visual space model has been integrated into MetaFormer for image restoration of brain Magnetic Resonance Imaging (MRI) and abdominal Computer Tomography (CT) [35]. For pancreas tissue detection, deformable convolutions have been employed as token mixers, dynamically adapting the receptive field [36]. To the best of our knowledge, this is the first comprehensive study of different standard token mixers in the MetaFormer on medical image analysis.
@@ 2 Methods
In a typical medical image processing task, we are given an input image $\mathbf{I}\in\mathbb{R}^{C\times H\times W}$ and aim to learn a function $f(\mathbf{I})$ that projects the image into a predefined solution space. In image classification, a vector in this solution space is described as a global vector $\mathbf{s}\in\mathbb{R}^{|\Omega|}$ holding the likelihood $p(\mathbf{I}|\omega_{n})$ for each class $\omega\in\Omega$. For semantic segmentation, the classification vector is defined for every pixel $\mathbf{p}\in\mathbf{I}$. Thus, we obtain a segmentation mask $\mathbf{S}\in\mathbb{R}^{|\Omega|\times H\times W}$ holding the likelihoods for every pixel $p(\mathbf{p}|\omega)$ to belong to class $\omega$. When approximating the optimal $f$ a given task with deep learning, we used to implement $f=f_{0}\circ f_{1}\circ\dots\circ f_{L}$ as a nested chain of functions $f_{l}$, where each $\{f_{l}\}_{l=0}^{L}$ represent a layer in a neural network. In the past ten years, the favoured operator in image processing for $f_{l}$ has been convolution [37, 38, 19]. After their initial success in natural language processing tasks, Transformers with their self-attention operator have also emerged as an alternative to convolutional networks [7, 8, 9]. However, while ViTs achieves state-of-the-art performance in many image tasks, it remains unclear, which architectural components drive their success. To address this, MetaFormer [23, 24] generalizes the Transformer architecture by replacing self-attention with operations better suited to vision.
@@ 2.1 MetaFormer Architecture
The MetaFormer [23] block represents a generalized version of the Transformer in that it defines token mixing as a variable operation, of which self-attention is only one of many possible operations. The other structures of the Transformers (e.g. channel MLP or residual connections) have been proven to be also crucial for computer vision [23] and are therefore left untouched. Fig. 1(b) shows the structure of the MetaFormer’s block. Within its multi-stage architecture, each employing multiple MetaFormer blocks, the MetaFormer adopts the hierarchical approach of CNNs [38, 37] of halving the spatial resolution and doubling the feature channels at the beginning of each stage using a convolution layer (see Fig. 1(c)). The first stage is an exception, where the downsample factor is four. The number of blocks per stage is a hyperparameter to modify model capacity, similar to the ResNet [38]. With this, the MetaFormer enables feature extraction on different resolution scales, which stands in contrast to the original ViT[7], where the image is patchified and then fed as a sequence of patches through the network without deviating from the original resolution.
(a) FLOPs of the token mixers across the four MetaFormer stages (stage index: $C\times N$) for an input of size $3\times 768^{2}$.
(b) MetaFormer block with different token mixer choices
(c) Extension of MetaFormer for global prediction (classification) by collapsing the final feature map to a vector and dense prediction (segmentation) employing the SegFormer decoder [8]. $|\Omega|$ describes the number of classes and $C=256$ represents the latent dimension of the decoder.
Figure 1: An overview of the MetaFormer’s architecture extended for global and dense prediction.
@@ 2.2 Different Token Mixers
Following MetaFormer [23, 24], we investigate commonly used operations as token mixers denoted by $t(x)$. We compare identity, average pooling, convolution, grouped convolution, self-attention, and local self-attention, all defined in the following paragraphs. The choice of token mixer introduces fundamental differences in properties such as the number of trainable parameters and runtime complexity. Moreover, each token mixer carries distinct inductive biases regarding locality, making them more appropriate for particular imaging domains (e.g., those requiring strong local interactions, long-range context, or texture-based representations).
The computational complexity and trainable parameter count of the investigated token mixers are summarized in Tab. 1We state the complexity in terms of the input size: number of channels $C$, height $H$, and width $W$ (for shorter notation we define $N=HW$). For local operators, we define their kernel or pooling size $K$, where we only consider kernels of size $(K\times K)$ containing $K^{2}$ elements, since it is the most common choice. Note that a MetaFormer block does not change the spatial resolution or number of channels and, therefore, we use the same variables for in- and output sizes.
To define the following operations, we introduce the indices $c,h,w\in\mathbb{N}$ iterating over the $C$ channels, the height $H$ and width $W$. We define the pixel value at the index $(c,h,w)$ as $\mathbf{I}(c,h,w)=p\in\mathbb{R}$. Further, let $\mathcal{N}$ define a function that returns the neighbouring pixels of a given location:
$\mathcal{N}_{K}(h,w)=\left\{i,j\in\mathbb{N}_{0}:\left(|i-h|<\frac{K}{2}\right)\land\left(|j-w|<\frac{K}{2}\right)\right\},$ (1)
where $K$ describes the kernel size in height and width, since we limit ourselves to symmetrical kernels.

### v2

# Shaken or Stirred? An Analysis of MetaFormer’s Token Mixing for Medical Imaging

@@ Abstract
The generalization of the Transformer architecture via MetaFormer has reshaped our understanding of its success in computer vision. By replacing self-attention with simpler token mixers, MetaFormer provides strong baselines for vision tasks. However, while extensively studied on natural image datasets, its use in medical imaging remains scarce, and existing works rarely compare different token mixers, potentially overlooking more suitable designs choices. In this work, we present the first comprehensive study of token mixers for medical imaging. We systematically analyze pooling-, convolution-, and attention-based token mixers within the MetaFormer architecture on image classification (global prediction task) and semantic segmentation (dense prediction task). Our evaluation spans eight datasets covering diverse modalities and common challenges in the medical domain. Given the prevalence of pretraining from natural images to mitigate medical data scarcity, we also examine transferring pretrained weights to new token mixers. Our results show that, for classification, low-complexity token mixers (e.g. grouped convolution or pooling) are sufficient, aligning with findings on natural images. Pretrained weights remain useful despite the domain gap introduced by the new token mixer. For segmentation, we find that the local inductive bias of convolutional token mixers is essential. Grouped convolutions emerge as the preferred choice, as they reduce runtime and parameter count compared to standard convolutions, while the MetaFormer’s channel-MLPs already provide the necessary cross-channel interactions. Our code is available at https://github.com/multimodallearning/MetaFormerMedImaging/tree/clean_code.
@@
doi:
10.59275/j.melba.2024-AAAA
volume: 3
CNN
Convolutional Neural Network
NLP
Natural Language Processing
ViT
Vision Transformer
MLP
Multi-Layer Perceptron
SAR
Synthetic Aperture Radar
CT
Computer Tomography
MRI
Magnetic Resonance Imaging
FLOPs
Floating Point Operations
AUC
Area under the Receiver Operating Characteristic curve
DSC
Dice Similarity Score
DETR
Detection Transformer
GAP
Global Average Pooling
@@ 1 Introduction
Over the last ten years, Deep Learning and especially Convolutional Neural Networks have set the state-of-the-art in various tasks of medical image processing like classification (Anwar et al., 2018), segmentation (Isensee et al., 2021) and detection (Fernando et al., 2021). Following its success in Natural Language Processing (NLP(Devlin et al., 2019; Radford et al., 2019), self-attention (Vaswani et al., 2017) has soon emerged as an alternative to convolution. The Transformer architecture has been successfully adapted to image classification (Dosovitskiy et al., 2021), segmentation (Xie et al., 2021) and detection (Carion et al., 2020). Further adaptation of the Transformer architecture targeted specific requirements for vision tasks. This includes for example a hierarchical pyramid-like architecture like it can be found in CNNs (Wang et al., 2021; Liu et al., 2021), the limitation of its receptive field to the direct neighbourhood for detection (Zhu et al., 2020), or tackling characteristics of self-attention that hinder the usage of Vision Transformers as the backbone for further vision tasks (Darcet et al., 2023).
@@
The property of self-attention having a global receptive field is a doubled-edged sword in image processing where large input sizes are common. On the one hand, the number of trainable parameters of the self-attention operator is independent of the input size, and it enables capturing of large context. On the other hand, since every pixel attends to each other pixel in an input, its computational complexity grows quadratically with the resolution of 2D images, making it inefficient for high resolutions. In contrast, convolution limits its receptive field to the direct neighbourhood of the pixel and thus provides a strong local bias and translation-equivariance, two properties fitting the domain of images. Combining both paradigms, many works have also limited the receptive field of self-attention, introducing local self-attention. Here, the implementations can be divided into two groups, splitting the input into patches (Vaswani et al., 2021) or limiting the influence of pixels outside the direct neighbourhood by applying a (soft) weighting during the self-attention operation (d’Ascoli et al., 2021). For both local and global attention, implementations were introduced (Pan et al., 2023; Dao et al., 2022). Further, recent work investigates into the principle of filter separation, bringing state space models to computer vision within the MAMBA architecture (Zhu et al., 2024). Lately, frameworks like PyTorch’s flexattention enables the easy implementation of various attention patterns.
@@
There also exist efforts (e.g. classification (Liu et al., 2022) and segmentation (Lee et al., 2022; Roy et al., 2023)) to enable large receptive fields for CNNs without equipping the model with too many learnable parameters.
@@
While Tolstikhin et al. (2021) demonstrate that replacing the self-attention within the ViT with Multi-Layer Perceptron (MLP) yields comparable performance (Dosovitskiy et al., 2021; Vaswani et al., 2021), the MetaFormer framework (Yu et al., 2022) takes this a step further by abstracting the general transformer architecture into two parts: The token mixer (e.g. self-attention) learns to aggregate spatial information and the channel-MLP extracts new features with a point-wise operation. Within this MetaFormer perspective, using grouped convolution as the token mixer naturally draws a parallel to ConvNeXt (Liu et al., 2022), where depthwise convolution fulfils an analogous role in aggregating spatial information. However, using pooling as a simple, non-parameterised token mixer, the MetaFormer outperforms ViT on ImageNet. This implies that the channel-MLPs provide the main contribution to the transformer’s capabilities in computer vision, rather than the token mixer. In their following work, Yu et al. (2023) provide strong results (ImageNet top-1 accuracy of over $80\,\%$) using no token mixing at all. Still, equipping the token mixers with some learnable parameters was found beneficial (see Yu et al. (2022), Tab. 5).
@@
The general formulation of the MetaFormer easily allows for the deployment of hybrid architectures, combining different token mixers (e.g. using pooling as an efficient token mixer in the first stages with high-resolution images and global self-attention in latter ones). Since then, MetaFormer has been implemented in various applications and extended with specialised token mixers (see Sec. 2).
@@ 1.1 Contribution
Most existing applications of the MetaFormer lie outside the medical domain, and those within typically employ it without comparing different token mixers – potentially overlooking more suitable choices. This work presents the first comprehensive study of token mixers for medical imaging, analysing their strengths, weaknesses, and complexity in terms of trainable parameters and runtime. We systematically compare six token mixers, varying the kernel size for kernel-based mixers such as convolution and pooling. Our evaluation covers eight datasets, including seven from the medical domain and one natural image benchmark. We investigate two different tasks: classification as a global prediction task (five datasets) and semantic segmentation as a dense prediction task (three datasets, covering small, medium, and large resolutions). The datasets represent diverse modalities, including histology, microscopy, and radiography. With this, we chose to expose the different token mixers to key medical imaging challenges such as large homogeneous regions in abdominal CT and long-range context requirements in pathology. We summarise our key findings as follows.
  • Datasets containing a single dominant object in the image favour more locality and thus smaller kernel sizes. Conversely, datasets with a larger region of interest tend to profit from incorporating a broader spatial context through larger kernels.
  • For classification, the MetaFormer trained from scratch with no token mixing already yields very competitive performance. The results are best when using low-complexity token mixers.
  • When pretrained weights are available, exchanging token mixers introduces a domain gap. But pretrained weights of global attention can still be successfully utilized in local attention. This provides a more efficient alternative to global attention when dealing with large input resolutions.
@@
2 Related Works
With the right choice of token mixer, the MetaFormer can be a light-weighted model, enabling its use for real-time applications like ship signature classification in Synthetic Aperture Radar (SAR) (Zhu et al., 2023) or in autonomous driving (Shun et al., 2024), which deploys additional token pruning for reduced runtime complexity. For super resolution of meteorological satellite cloud images, Wang et al. (2023) introduce a new token mixer combining gated convolution and channel-space attention. With the extension of temporal pooling, the MetaFormer has been applied on time series, capturing the structural and temporal relationship in human motion prediction for human-robot interaction (Xu et al., 2023). A novel idea of token mixing moves this operation to the Fourier spectrum, enabling efficient global token mixing without limited receptive field, but adding the cost of the Fourier transformation and its inverse. With balancing the low- and high-pass characteristics via spectral pooling aggregation (Yun et al., 2023), the performance on ImageNet is improved by $3\,\%$ accuracy compared to the PoolFormer (MetaFormer with average pooling as token mixer) (Yu et al., 2022). In Tatsunami and Taki (2024), this approach is extended, so coefficients of a dynamic filter bank are estimated by an MLP based on the input. While the original work (Yu et al., 2022) employs the MetaFormer itself as a backbone for semantic segmentation, MetaSeg (Kang et al., 2024) extends its principles to a decoder and introduces channel reduction attention for gathering global context within the decoder.
In the medical domain, MetaFormer mostly gets adapted for semantic segmentation. For polyp segmentation, MetaFormer has been paired with a parallel path comprising convolutional units with residual and dilated paths to compensate for MetaFormer’s weak low-level features (Lee and Yoo, 2024). Another work on polyp segmentation (Trinh et al., 2023) utilizes the MetaFormer as the encoder in a U-Net-like architecture and extend the decoder’s dilated convolutions with learnable multiscale upsampling. In volumetric brain tumour segmentation, MetaFormer has been used as an auxiliary decoder, combining local–global pooling with depthwise convolution as the token mixer. To improve efficiency, it has been adapted with involution and a squeeze-and-excitation module in place of channel-MLP (Dwika Hefni Al-Fahsi et al., 2024). Similarly, MAMBA’s structured visual space model has been integrated into MetaFormer for image restoration of brain Magnetic Resonance Imaging (MRI) and abdominal Computer Tomography (CT) (Chen et al., 2024). For pancreas tissue detection, deformable convolutions have been employed as token mixers, dynamically adapting the receptive field (Kurami et al., 2024). To the best of our knowledge, this is the first comprehensive study of different standard token mixers in the MetaFormer on medical image analysis.
@@ 3 Methods
Figure 1: Schema of the MetaFormer block illustrating different token mixer options employed across stages of the MetaFormer architecture (cf. Fig. 2(b)).
((a)) FLOPs of the token mixers across the four MetaFormer stages (stage index: $C\times N$) for an input of size $3\times 768^{2}$.
((b)) Extension of MetaFormer for classification by collapsing the final feature map to a vector and segmentation employing the SegFormer decoder Xie et al. (2021) with number of classes $|\Omega|$ and latent dimension $C=256$ for segmentation.
Figure 2: Visualization of FLOPs for different token mixers (2(a)) and schema of MetaFormer’s architecture (2(b)).
In a typical medical image processing task, we are given an input image $\mathbf{I}\in\mathbb{R}^{C\times H\times W}$ and aim to learn a function $f(\mathbf{I})$ that projects the image into a predefined solution space. In image classification, a vector in this solution space is described as a global vector $\mathbf{s}\in\mathbb{R}^{|\Omega|}$ holding the likelihood $p(\mathbf{I}|\omega_{n})$ for each class $\omega\in\Omega$. For semantic segmentation, the classification vector is defined for every pixel $\mathbf{p}\in\mathbf{I}$. Thus, we obtain a segmentation mask $\mathbf{S}\in\mathbb{R}^{|\Omega|\times H\times W}$ holding the likelihoods for every pixel $p(\mathbf{p}|\omega)$ to belong to class $\omega$. When approximating the optimal $f$ in a given task with deep learning, we used to implement $f=f_{0}\circ f_{1}\circ\dots\circ f_{L}$ as a nested chain of functions $f_{l}$, where each $\{f_{l}\}_{l=0}^{L}$ represent a layer in a neural network. In the past ten years, the favoured operator in image processing for $f_{l}$ has been convolution (Krizhevsky et al., 2012; He et al., 2016; Liu et al., 2022). After their initial success in natural language processing tasks, Transformers with their self-attention operator have also emerged as an alternative to convolutional networks (Dosovitskiy et al., 2021; Xie et al., 2021; Carion et al., 2020). However, while ViTs achieve state-of-the-art performance in many image tasks, it remains unclear, which architectural components drive their success. To address this, MetaFormer (Yu et al., 2022, 2023) generalizes the Transformer architecture by replacing self-attention with operations better suited to vision.
@@ 3.1 MetaFormer Architecture
The MetaFormer (Yu et al., 2022) block represents a generalized version of the Transformer in that it defines token mixing as a variable operation, of which self-attention is only one of many possible operations. The other structures of the Transformers (e.g. channel MLP or residual connections) have been proven to be also crucial for computer vision (Yu et al., 2022) and are therefore left untouched. Fig. 1 shows the structure of the MetaFormer’s block. Within its multi-stage architecture, each employing multiple MetaFormer blocks, the MetaFormer adopts the hierarchical approach of CNNs (He et al., 2016; Krizhevsky et al., 2012) of halving the spatial resolution and doubling the feature channels at the beginning of each stage using a convolution layer (see Fig. 2(b)). The first stage is an exception with a downsampling factor of four. The number of blocks per stage is a hyperparameter to modify model capacity, similar to the ResNet (He et al., 2016). With this, the MetaFormer enables feature extraction on different resolution scales, which stands in contrast to the original ViT (Dosovitskiy et al., 2021), where the image is patchified and then fed as a sequence of patches through the network without deviating from the original resolution.
@@ 3.2 Different Token Mixers
Following MetaFormer (Yu et al., 2022, 2023), we investigate commonly used operations as token mixers denoted by $t(x)$. We compare identity, average pooling, convolution, grouped convolution, self-attention, and local self-attention, all defined in the following paragraphs. The choice of token mixer introduces fundamental differences in properties such as the number of trainable parameters and runtime complexity. Moreover, each token mixer carries distinct inductive biases regarding locality, making them more appropriate for particular imaging domains (e.g., those requiring strong local interactions, long-range context, or texture-based representations).
The computational complexity and trainable parameter count of the investigated token mixers including the channel-MLP of the MetaFormer block are summarized in Tab. 1. We state the complexity in terms of the input size: number of channels $C$, height $H$, and width $W$ (for shorter notation we define $N=HW$). For local operators, we define their kernel or pooling size $K$, where we only consider kernels of size $(K\times K)$ containing $K^{2}$ elements, since it is the most common choice. Note that a MetaFormer block does not change the spatial resolution or number of channels and, therefore, we use the same variables for in- and output sizes.
To define the following operations, we introduce the indices $c,h,w\in\mathbb{N}$ iterating over the $C$ channels, the height $H$ and width $W$. We define the pixel value at the index $(c,h,w)$ as $\mathbf{I}(c,h,w)=p\in\mathbb{R}$.